# SpecReason: Fast and Accurate Inference-Time Compute via Speculative Reasoning

### v1

@@ 4.2 Hierarchical Speculation across Semantic Similarity and Token Equivalence
At a high level, SpecReason’s speculative reasoning resembles the philosophy behind traditional speculative decoding, but differs in two important ways. First, speculative decoding guarantees token-level equivalence between draft and verified outputs, making it a form of exact acceleration. In contrast, SpecReason targets semantic-level similarity, accepting steps that carry the same insight even if phrased differently. Second, speculative decoding is typically applied to output generation tasks (e.g., text continuation or translation), where the fidelity of each token matters. SpecReason, on the other hand, is designed specifically for internal thinking tokens in reasoning tasks, where intermediate steps are approximate and interchangeable as long as they preserve the logical progression of thought.
@@ 5.1 Setup
Models. In our main results, we use QwQ-32B qwq (2025) as the base model and a distilled version of DeepSeek-R1, R1-1.5B Guo et al. (2025), as the small model – both of which are fine-tuned on Qwen-2.5 Yang et al. (2024) and embed the capability of reasoning with long CoTs. We evaluate an additional base model with a different size and architecture, R1-70B Guo et al. (2025), a distilled version of DeepSeek-R1 onto Llama3.3-70B Grattafiori et al. (2024), in §5.4.
Datasets. We evaluate SpecReason on three diverse reasoning benchmarks: AIME aim (2025) for high-school competition-level mathematical problems, MATH500 Hendrycks et al. (2021) for high-school competition-level mathematical problems sampled from AMC 10, AMC 12, and AIME, and GPQA Diamond Rein et al. (2024) for graduate-level questions in general domains like biology, physics, and chemistry. The accuracy metric we evaluate on is pass@1. Similar to prior work Guo et al. (2025), we set k=16 when calculating pass@1 – i.e., we generate 16 responses with temperature=0.6 for every query and calculate the average accuracy – and set the token budget to be 8192 tokens to ensure an apples-to-apples comparison between baselines.
Baselines. We run vanilla inference using the small model and the base model as the latency baseline and accuracy baseline, respectively. Aside from SpecReason, we also run speculative decoding (“SpecDecode”) with R1-1.5B as the draft model, speculating five tokens at a time. To demonstrate SpecReason’s compatibility with speculative decoding, we also run a “SpecReason+Decode” baseline that employs the hierarchical speculation described in §4.2.
@@
Figure 3: Comparison of the accuracy and latency of different schemes. SpecReason significantly reduces latency while improving accuracy over vanilla inference, and is also complementary with speculative decoding.
Figure 4: SpecReason allows trading off latency for accuracy via adjusting the acceptance threshold (from left to right, the thresholds are: 3, 5, and 7 out of 9).
@@ 5.2 Main Results
We compare SpecReason against baseline methods in Fig. 3. Across the three datasets—MATH, AIME, and GPQA—SpecReason achieves a 2.5$\times$, 1.9$\times$, and 1.5$\times$ reduction in latency, respectively, compared to vanilla inference with the base model. Alongside these gains, it also yields modest accuracy improvements of 1.0%, 3.1%, and 9.9% compared to the base model. We attribute these accuracy gains to SpecReason’s explicit judgment and scoring mechanism at each reasoning step, which augments the model’s internal self-reflection with more structured assessment.
When compared with speculative decoding, SpecReason lies on the Pareto frontier of the accuracy-latency tradeoff. More importantly, combining SpecReason with speculative decoding (SpecReason+Decode) results in further latency reductions of 44.2%, 33.8%, and 19.4% over speculative decoding alone. The most significant performance gains for SpecReason occur on the MATH dataset, where both models achieve relatively high accuracies and the capability gap between the small and base models is the narrowest. This makes intermediate steps easier for the small model to speculate correctly, increasing the acceptance rate of speculated steps and thereby lowering end-to-end latency.
Finally, when comparing SpecReason+Decode with SpecReason, SpecReason+Decode reduces latency by 1.7$\times$, 1.9$\times$, and 1.7$\times$, demonstrating the difference in ease of speculation across varying tasks. On these three datasets, the ratio of steps carried out by small models in SpecReason is 80.0%, 40.8%, and 44.4%, respectively.
@@ 5.3 Controlling the Accuracy-Latency Tradeoff
In Fig. 4, we illustrate how SpecReason enables flexible control over the accuracy-latency tradeoff, using a representative, randomly selected subdataset from the full datasets in §5.4 for ease of evaluation. During the base model’s evaluation of each reasoning step, we vary the acceptance threshold for the utility score between 3, 5, and 7, and report the resulting accuracy and latency.
On the MATH subdataset, increasing the acceptance threshold from 3 to 7 results in fewer speculative steps from the small model being accepted. This leads to a latency increase from 35.7s to 69.2s, while accuracy improves from 59.4% to 63.7%, due to tighter control over the approximation level of intermediate reasoning steps. Notably, the gap between SpecReason+Decode and SpecReason widens from 8.1s to 28.8s, since more reasoning steps are delegated to the base model and SpecReason+Decode reduces only the base model’s decoding time compared to SpecReason.
A similar trend is observed on the AIME subdataset: as the acceptance threshold increases from 3 to 7, latency grows from 109.4s to 261.9s, and accuracy improves from 22.9% to 40.5%. However, the accuracy degrades less gracefully as the threshold is relaxed compared to the MATH subdataset. This is because the small model exhibits a larger performance gap relative to the base model on AIME, making aggressive acceptance of its speculative steps more costly in terms of accuracy.
5.4 Base Models of Varying Sizes and Architectures
@@
Figure 6: The utility scores in SpecReason closely reflects the quality score judgements from a process reward model. $x$ on the x-axis denotes PRM scores in the range $[x,x+0.1)$.
@@
5.5 Base Model’s Judgement Capability
The base model’s ability to assess the quality of intermediate reasoning steps is a crucial cornerstone of SpecReason’s performance. In this experiment, we compare the scores generated by a process reward model (PRM) – which assigns a reward score to each step within the solution to a math problem – with those given by the QwQ-32B base model on the AIME dataset. Specifically, we use Math-Shepherd Wang et al. (2023), a PRM trained via reinforcement learning from the Mistral-7B base model on math problems, to score each speculated step produced by the R1-1.5B small model.
6 Related Work
There has been some early work aimed at reducing the inference latency of LRMs. Sky-T1-Flash Team (2025) addresses the problem of overthinking by fine-tuning models to curb unnecessary reasoning, thereby reducing the length of reasoning chains and, consequently, latency. While this approach reduces computation by curtailing reasoning depth, it does not fully leverage the potential of doing approximations on easier intermediate reasoning steps for efficiency.
Dynasor-CoT Fu et al. (2024b, 2025) takes a different approach by probing intermediate model confidence and terminating the reasoning process early when the model exhibits sufficient confidence in its current output. This early-exit strategy helps mitigate latency and is complementary to SpecReason, as both aim to reduce unnecessary reasoning steps.
7 Conclusion
In this work, we introduce SpecReason, a novel approach that accelerates LRM inference by leveraging speculative reasoning. By offloading simpler intermediate reasoning steps to a smaller, lightweight model and reserving the base model for assessment, SpecReason significantly reduces inference latency while maintaining or even improving accuracy. Our results demonstrate that SpecReason achieves a 1.5–2.5× speedup over traditional LRM inference, with accuracy improvements ranging from 1.0–9.9%. Additionally, when combined with speculative decoding, SpecReason further reduces latency by 19.4–44.2%, highlighting the complementary nature of these optimizations. We believe this work opens up new angles for efficient LRM inference acceleration, making it especially valuable for scenarios that demand both high accuracy and low latency.
Acknowledgments and Disclosure of Funding

### v2

@@ 4.2 Hierarchical Speculation across Semantic Similarity and Token Equivalence
At a high level, SpecReason’s speculative reasoning resembles the philosophy behind traditional speculative decoding, but differs in two important ways. First, speculative decoding guarantees token-level equivalence between draft and verified outputs, making it a form of exact acceleration. In contrast, SpecReason targets semantic-level similarity, accepting steps that carry the same insight even if phrased differently, and exposes knobs to control the exactness of reasoning approximations. Second, speculative decoding is typically applied to output generation tasks (e.g., text continuation or translation), where the fidelity of each token matters. SpecReason, on the other hand, is designed specifically for internal thinking tokens in reasoning tasks, where intermediate steps are approximate and interchangeable as long as they preserve the logical progression of thought.
@@ 5 Evaluation
The overview of our evaluation results include:
  • $\bullet$
    Reducing end-to-end latency. Because many intermediate steps are easier than end-to-end reasoning, many (up to 80%) of the speculated steps are accepted. SpecReason achieves a 1.43.0×1.4-3.0$\times$1.4 - 3.0 × speedup over vanilla LRM inference. Additionally, when combined with speculative decoding, SpecReason further reduces latency by $8.8-58.0\%$ over speculative decoding alone, highlighting the complementary nature of these optimizations.
  • $\bullet$
    Improving token-budget-aware accuracy. Beyond latency reduction, SpecReason also improves accuracy over the base model by $0.4-9.0\%$ under the same token budget. We empirically find that small, lightweight models typically have shorter output sequence lengths – meaning, they need fewer thinking tokens before deriving an answer. Thus, by accepting many small model’s speculated reasoning steps, SpecReason reduces the token consumption compared to the base model’s vanilla inference. When the token budget is low – a common setup to curb inference cost and latency – SpecReason helps improve accuracy as the base model would need more tokens to get to an answer (Fig. 4).
@@ 5.1 Setup
Models. In our main results, we use two base models: QwQ-32B (qwq, 2025) and Skywork-OR1-Preview-32B (sky, 2025). We also use two different small models for speculation: DeepSeek-R1-1.5B (Guo et al., 2025) and Zyphra’s ZR1-1.5B (zyp, 2025) – both of which are based on Qwen-2.5 (Yang et al., 2024) and embed the capability of reasoning with long CoTs – and evaluate all four different model combinations. We evaluate an additional base model with a different size and architecture, R1-70B (Guo et al., 2025), a distilled version of DeepSeek-R1 onto Llama3.3-70B (Grattafiori et al., 2024), in §A.1.
Datasets. We evaluate SpecReason on three diverse reasoning benchmarks: AIME (aim, 2025) for high-school competition-level mathematical problems, MATH500 (Hendrycks et al., 2021) for high-school competition-level mathematical problems sampled from AMC 10, AMC 12, and AIME, and GPQA Diamond (Rein et al., 2024) for graduate-level questions in general domains like biology, physics, and chemistry. The accuracy metric we evaluate on is pass@1. Similar to prior work (Guo et al., 2025), we set k=16 when calculating pass@1 – i.e., we generate 16 responses with temperature=0.6 for every query and calculate the average accuracy – and set the token budget to be 8192 tokens to ensure an apples-to-apples comparison between baselines.
Baselines. We run vanilla inference using the small and base models as the latency and accuracy baseline, respectively. Aside from SpecReason, we also run speculative decoding (“SpecDecode”) with the smaller model as the draft model, speculating five tokens at a time. To demonstrate SpecReason’s compatibility with speculative decoding, we also run a “SpecReason+Decode” baseline that employs the hierarchical speculation described in §4.2.
@@
(a) QwQ-32B + R1-1.5B
(b) QwQ-32B + Zyphra-1.5B
(c) Skywork-Preview-32B + R1-1.5B
(d) Skywork-Preview-32B + Zyphra-1.5B
Figure 3: Comparison of the accuracy and latency of different schemes on different model combinations. SpecReason significantly reduces latency while improving accuracy over vanilla inference. When combined with speculative decoding, SpecReason outperforms speculative decoding in both latency and accuracy on all datasets and model combinations.
(a) Output length comparison. SpecReason reduces the token consumption needed to answer queries by adopting speculated steps from small models that are less verbose.
(b) [AIME] Accuracy gap under different token budgets.
Figure 4: [QwQ-32B + Zyphra-1.5B] Intuition behind SpecReason’s accuracy improvement. See Fig. 9 in §A for the full set of results.
Figure 5: [QwQ-32B + R1-1.5B] SpecReason allows trading off latency for accuracy via adjusting the acceptance threshold (from left to right, the thresholds are: 3, 5, 7, and 9 out of 9).
@@ 5.2 Main Results
We compare SpecReason against baseline methods in Fig. 3. Across the four model combinations, SpecReason achieves a 1.5$\times$–2.5$\times$, 1.6$\times$–3.0$\times$, 1.4$\times$–2.5$\times$, 1.7$\times$–2.4$\times$ reduction in latency, respectively, compared to vanilla inference with the base model.
Accuracy improvement. Alongside these efficiency gains, SpecReason also yields modest accuracy improvements of 1.3%–3.6%, 4.0%–9.0%, 0.4%–1.7%, and 1.4%–5.0% compared to the base model. The key reason behind this accuracy improvement is the reduction in token consumption required for reasoning. In Fig. 4, we focus on the model combination with the highest overall accuracy improvement, QwQ-32B + Zyphra-1.5B, and compare the average number of thinking tokens needed to derive an answer between the base model, the small model, and SpecReason. As seen in Fig. 4(a), the small model is generally less verbose than the base model, and because SpecReason adopts many speculated steps from the small model, its token consumption is also reduced by 1.2$\times$–2.0$\times$. We also focus on the AIME dataset and vary the token budget to study its effect on the difference in accuracy between SpecReason and the base model in Fig. 4(b). The effect of token reduction on accuracy is the most significant for tighter output token budgets (16.2% at 4096 tokens) but shrinks as the base model is allowed to generate more thinking tokens (4.7% at 8192 tokens). We also attribute these accuracy gains to SpecReason’s explicit judgment and scoring mechanism at each reasoning step, which augments the model’s internal self-reflection with more structured assessment.
When compared with speculative decoding, SpecReason lies on the Pareto frontier of the accuracy-latency tradeoff. More importantly, combining SpecReason with speculative decoding (SpecReason+Decode) results in further latency reductions of 19.4%–44.2%, 30.8%–58.0%, 8.8%–52.2%, and 25.1%–51.8% over speculative decoding alone. The most significant performance gains for SpecReason when the base model is QwQ-32B occur on the MATH dataset, where both models achieve relatively high accuracies and the capability gap between the small and base models is the narrowest. This makes intermediate steps easier for the small model to speculate correctly, increasing the acceptance rate of speculated steps and thereby lowering end-to-end latency. In comparison, Skywork-Preview-32B is slightly inferior at instruction following, so SpecReason has to adopt a higher threshold to avoid an accuracy loss, reducing SpecReason’s latency wins.
Finally, when comparing SpecReason+Decode with SpecReason, SpecReason+Decode reduces latency by 1.7$\times$–1.9$\times$, 1.7$\times$–1.8$\times$, 1.6$\times$–2.2$\times$, and 1.6$\times$–2.1$\times$, demonstrating the difference in ease of speculation across varying tasks. On these three datasets, the ratio of steps carried out by small models in SpecReason is 38.1%–80.0%, 36.5%–71.3%, 39.3%–70.2%, and 41.4%–66.6%, respectively.
@@ 5.3 Controlling the Accuracy-Latency Tradeoff
In Fig. 5, we illustrate how SpecReason enables flexible control over the accuracy-latency tradeoff, using a representative, randomly selected subdataset from the full datasets in §5.2 on QwQ-32B + R1-1.5B for ease of evaluation. During the base model’s evaluation of each reasoning step, we vary the acceptance threshold for the utility score between 3, 5, 7, and 9, and report the resulting accuracy and latency.
On the MATH subdataset, increasing the acceptance threshold from 3 to 7 results in fewer speculative steps from the small model being accepted. This leads to a latency increase from 35.7s to 69.2s, while accuracy improves from 59.4% to 63.7%, due to tighter control over the approximation level of intermediate reasoning steps. Notably, the gap between SpecReason+Decode and SpecReason widens from 8.1s to 28.8s, since more reasoning steps are delegated to the base model, and SpecReason+Decode reduces only the base model’s decoding time compared to SpecReason.
A similar trend is observed on the AIME and GPQA subdatasets: as the acceptance threshold increases from 3 to 7, latency grows from 109.4s to 261.9s and from 72.7s to 223.0s, and accuracy improves from 22.3% to 39.3% and from 33.1% to 50.7%. However, the accuracy degrades less gracefully as the threshold is relaxed compared to the MATH subdataset. This is because the small model exhibits a larger performance gap relative to the base model on AIME and GPQA, making aggressive acceptance of its speculative steps more costly in terms of accuracy.
In Fig. 7, we also study the effect of the alternative knob, forcing the first $n$ reasoning steps to be decoded by the base model, on the accuracy-latency tradeoff. As we change $n$ from 0 to 10, 20, 30, and 40, SpecReason’s accuracy increases from 33.2% to 37.3% while the latency increases from 270.4s to 292.6s, showcasing an alternative approach to improve accuracy with a slight increase in latency.
Figure 6: Effect of the alternative knob: forcing the first $n$ steps for base model decoding.
Figure 7: The utility scores in SpecReason closely reflect the quality score judgements from a process reward model. $x$ on the x-axis denotes PRM scores in the range $[x,x+0.1)$.
5.4 Base Model’s Judgement Capability
The base model’s ability to assess the quality of intermediate reasoning steps is a crucial cornerstone of SpecReason’s performance. In this experiment, we compare the scores generated by a process reward model (PRM) – which assigns a reward score to each step within the solution to a math problem – with those given by the QwQ-32B base model on the AIME dataset. Specifically, we use Math-Shepherd (Wang et al., 2023), a PRM trained via reinforcement learning from the Mistral-7B base model on math problems, to score each speculated step produced by the R1-1.5B small model.
6 Conclusion
In this work, we introduce SpecReason, a novel approach that accelerates LRM inference by leveraging speculative reasoning. By offloading simpler intermediate reasoning steps to a smaller, lightweight model and reserving the base model for assessment, SpecReason significantly reduces inference latency while maintaining or even improving accuracy. Our results demonstrate that SpecReason achieves a 1.43.0×1.4-3.0$\times$1.4 - 3.0 × speedup over vanilla LRM inference, with accuracy improvements ranging from $0.4-9.0\%$. Additionally, when combined with speculative decoding, SpecReason further reduces latency by $8.8-58.0\%$, highlighting the complementary nature of these optimizations. We believe this work opens up new angles for efficient LRM inference acceleration, making it especially valuable for scenarios that demand both high accuracy and low latency.
Acknowledgments and Disclosure of Funding
Appendix A Appendix
A.1 Base Models of Varying Sizes and Architectures
@@
A.2 Intuition behind Accuracy Improvement
Figure 9: Intuition behind SpecReason’s accuracy improvement on all datasets and model combinations.
In Fig. 9, we evaluate the average thinking token count of SpecReason and two vanilla inference baselines on a wide range of datasets and model combinations. We observe that the small model is generally less verbose than the base model, and because SpecReason adopts many speculated steps from the small model, its token consumption is reduced by 1.01.3×1.0-1.3$\times$1.0 - 1.3 ×, 1.22.0×1.2-2.0$\times$1.2 - 2.0 ×, 1.01.8×1.0-1.8$\times$1.0 - 1.8 ×, and $\times1$ 2.3 × on the four model combinations, respectively.